If Herons formula is used to find the area of a triangle ABC having a : 3, B=5 and (:6
Answer #1

heron's formula for area of triangle is given by

Area=\sqrt{s(s-a)(s-b)(s-c)}

where s is the semi perimeter ,

s=\frac{a+b+c}{2}

here given triangle sides as

a=3

b=5

c=6

we get

s=\frac{3+5+6}{2}=\frac{14}{2}=7

we get area as

Area=\sqrt{7(7-3)(7-5)(7-6)}

  =\sqrt{7(4)(2)(1)}

=\sqrt{56}=\boldsymbol{2\sqrt{14} \ sq.unit}